TyDaGreatest:

Bethany wants to estimate the value of (4.296 times 10 Superscript 11) (1.8614 times 10 Superscript negative 14). Which statement about the estimate is true?

QuestionCoveBot:

Can you post the statements please?

TyDaGreatest:

The value will be greater than 1 because 4.296 times 10 Superscript 11 is a very large number and multiplication always increases the size of a number. The value will be greater than 1 because 4 times 2 = 8 and 8 is larger than 1. The value will be less than 1 because (4.296 times 10 Superscript 11) (1.8614 times 10 Superscript negative 14) will be a negative number. All negative numbers are less than one. The number will be less than 1 because when adding the exponents, 11 + (negative 14) = negative 3, and a number in scientific notation with a negative exponent is less than 1.

QuestionCoveBot:

@snowflake0531

Extrinix:

\((4.296 \times 10 ^\text{11})(1.8614 \times 10 ^\text{-14})\) Now, to make it easier i suggest you solve it, then look at the answers and see which of those fit the answer best.

AZ:

You should know this property \( a^b \times a^c = a^{b+c}\) so what would \( 10^{11} \times 10^{-14}\) be ? and then another property you should know is \( x^{-y} = \dfrac{1}{x^y}\) so \(\bf for ~example\) if you had \(2^{-7}\) that would be \(\dfrac{1}{2^7}\) which is equal to \(\dfrac{1}{128} =0.0078125\) So having a negative exponent doesn't make the result negative, it just makes it a number that's very small. Something greater than 0 but less than 1 Does that help you estimate what the value is going to be for \((4.296 \times 10 ^{11})(1.8614 \times 10 ^{-14})\)